OpenStudy (anonymous):

prove sin(pi/2-x)=sin(pi/2+x)

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

I'd do it this way, using the sum and difference of sin expansions \[\sin(\frac{\pi}{2}- x) = \sin(\frac{\pi}{2})\cos(x) - \cos(\frac{\pi}{2})\sin(x)\] and \[\sin(\frac{\pi}{2} + x) = \sin(\frac{\pi}{2})\cos(x) + \cos(\frac{\pi}{2})\sin(x)\] simplify each... and see what you get...? you need to remember \[\sin(\frac{\pi}{2}) = 1 ~~~~~~~\cos(\frac{\pi}{2}) = 0\] hope it helps

OpenStudy (anonymous):

@ganeshie8

OpenStudy (xapproachesinfinity):

@campbell_st has given you what you need just simplify sin(pi/2) is what? cos(pi/2) is what?

OpenStudy (anonymous):

That was how i did, but it turns out its not the way to prove according to my professor

OpenStudy (anonymous):

and i think he is right because if u solve for the other part ull get a cosx

OpenStudy (xapproachesinfinity):

so what does you teachers wants? did he indicate what you should use? for example we can use cofunction of sin sin(pi/2-x)=cos(x) sin(pi/2+x)=sin(pi/2-(-x))=cos(-x)

OpenStudy (anonymous):

both side most be equal to each other

OpenStudy (xapproachesinfinity):

and since cosine is an even function so what would cos(-x) be

OpenStudy (xapproachesinfinity):

of course they should equal to each other

OpenStudy (anonymous):

cos(x)

OpenStudy (anonymous):

yes but but professor said not the way to solve this

OpenStudy (anonymous):

i have tried all possible methods that i know

OpenStudy (anonymous):

to find a way to prove that one side is equal the other but the answer isnt cosx

OpenStudy (anonymous):

the hall class did the way we just did and he said not the way to prove it

OpenStudy (anonymous):

whole

OpenStudy (anonymous):

well thanks

OpenStudy (anonymous):

you deserve a medal :D

OpenStudy (xapproachesinfinity):

idk what your prof want, there is a geometric proof but it is also using cos

ganeshie8 (ganeshie8):

so you want to take the expression on left hand side and deduce the right hand side, eh ?

OpenStudy (anonymous):

well if that will work would you mind walking me through it

ganeshie8 (ganeshie8):

without doing that cosx business

ganeshie8 (ganeshie8):

i see what ur professor wants, it looks tricky though

ganeshie8 (ganeshie8):

@Kainui

OpenStudy (anonymous):

for the past 2 days its the only thing in my mind lol

OpenStudy (kainui):

Well since sine is an odd function, \[\LARGE \sin(\theta)=-\sin(-\theta)\] and since sine is periodic, after half a period it's its own negative. \[\LARGE \sin(\theta)=-\sin(\theta + \pi)\] This should be all you need to find the answer.

OpenStudy (anonymous):

wow awesome

OpenStudy (anonymous):

i could never think of this lol just saying

OpenStudy (kainui):

I didn't invent it either, don't worry I just read about it and see that it makes sense and understand it and now it's one of my tools in solving problems =) You just have to look at the graph to see that it's true. \[\LARGE \sin(x)=-\sin(-x)\] Look at the graph of this, we can see that if we start putting in 1, 2, 3, etc... on the left side we get sine wiggling to the right. But if we put in 1, 2, 3, etc... on the right equation -sin(-x) then we get the values of sine as if you were moving to the left. Since those ones are upside down, we have to flip them with another negative sign.

OpenStudy (anonymous):

well thank you very much

OpenStudy (anonymous):

can you also send me link of where you read it.. plz

OpenStudy (kainui):

I don't have links to the inside of my brain. I don't know, I read and learned about this a long time ago, but you are free to ask me questions for the next 30 minutes and I'll help you before I go out partying for halloween =P

OpenStudy (anonymous):

hahaha :D i guess that will be it for today.. ill keep in touch :D thanks to you and @ganeshie8 and @xapproachesinfinity and @campbell_st

OpenStudy (anonymous):

you guys are the best 4real

OpenStudy (xapproachesinfinity):

Anytime!

ganeshie8 (ganeshie8):

watch first 4 minutes of this video to find out why your professor thinks ur proof is not good http://www.youtube.com/watch?v=NuGDkmwEObM

OpenStudy (xapproachesinfinity):

Eh I agree with the professor!